# BlendScape: Enabling Unified and Personalized Video-Conferencing Environments through Generative AI

### v1

@@ 5. Demonstration of Additional Scenarios
In the first part of our summative evaluation, we use scenario demonstration (Ledo et al., 2018) as a method for illustrating the range of examples that can be achieved with our system. Our review of prior work surfaced ten environment composition techniques for supporting distributed collaboration (Sec. 2.1), e.g., enabling proxemic metaphors, incorporating representations of the task space, or recording a history of collaborative activities. In this section, we describe how we used BlendScape (Ledo et al., 2018) to implement three distributed collaboration scenarios (Design Brainstorming, Remote Education, Storytime with Family) that incorporate seven of these techniques. Comparing our generated meeting spaces to prior video-conferencing environments, we analyze how BlendScape can be used to implement these techniques similarly to prior work or introduce new composition strategies.
@@ 5.1. Design Brainstorming
Recall our initial scenario from Sec. 3, where two interaction designers are collaborating remotely (Fig. 3). They use BlendScape to render their videos in a unified workspace that blends together their physical and digital task spaces, allowing them to refer to hand-drawn and digital sketches.
Implementation with BlendScape: We prototyped this scenario using three live video feeds: two videos of the designers captured through laptop webcams, and one external camera feed capturing a physical desk with notebooks and a digital sketch (Fig. 3b).
Comparison to Prior Work: Similar to MirrorBlender (Grønbæk et al., 2021), BlendScape environments afford layering users’ videos on top of shared content to enable more natural interactions with the task space (e.g., using deictic expressions, pointing to content via hand gestures). However, BlendScape approach of blending video feeds together into a seamless depiction of the physical and digital task spaces moves beyond “windowing” environments in Teams, Zoom, or video conferencing research systems (e.g.,(Grønbæk et al., 2021)), where elements of the task space are displayed in separate blocks. While further research is required to understand the impact of these blended spaces on collaboration workflows, prior empirical studies of video-conferencing (Tang et al., 2023) and mixed reality collaboration systems (Pejsa et al., 2016; Grønbæk et al., 2023b) suggest that depicting a spatial arrangement of users within a unified space can increase co-presence and improve conversation flows.
@@ 5.2. Remote Education
Next, to demonstrate how BlendScape environments could support proxemic and spatial metaphors to facilitate conversations (Hu et al., 2023a), we prototyped a scenario involving a remote university course.
Scenario Description: A professor uses BlendScape to facilitate activities in their literature seminar course. To establish the room layout for small group discussions, they upload an image prior of a room with several tables and restyle it as a “library” (Fig. 10a, c). Students split into groups by organizing themselves around the tables, leveraging proxemic metaphors (Fig. 10c). To structure a larger group discussion afterwards, the professor restyles an image of a conference room (Fig. 10b, d). To spotlight students who are speaking, they use scaling metaphors by rendering the students’ videos larger and at the head of the table.
Implementation with BlendScape: We used BlendScape’s image-to-image mode with a Meeting Theme prompt of “library” to transform the image priors. To prototype the proxemic and scaling metaphors, we used BlendScape’s direct manipulation techniques.
Comparison to Prior Work: In contrast with OpenMic (Hu et al., 2023a), which leverages abstract spatial metaphors to enable proxemic interactions (i.e., a Virtual Floor boundary), BlendScape’s environments introduce literal and concrete spatial landmarks for users to organize around (e.g., conference room table with chairs). This is similar to Gather’s approach, which uses walls to indicate meeting space boundaries and podiums to designate areas to broadcast audio. However, a key difference from Gather is that BlendScape’s environments are not pre-canned; unexpected spatial properties could encourage new collaborative behaviors (e.g., collaborators could be compelled to split into groups if the room geometry introduces barriers, as shown in Fig. 13). Future work is needed to study the impact of abstract vs. concrete and deterministic vs. stochastic spatial properties on distributed collaborators’ behavior.
@@ 5.3. Storytime with Family
Scenario Description: A grandmother uses BlendScape to craft a memorable storytelling experience with her granddaughter. First, with a prompt of “storytelling,” a castle appears in the blended space between their video backgrounds (Fig. 11b). As the grandmother reads a fairy tale, she restyles the scenes to set the theme of the meeting; the environment transforms into magic ballroom and then a mushroom forest (Fig. 11c, e). She and her granddaughter reposition their videos in the scene to enact the roles of different characters. A few weeks later, the granddaughter wants to write a sequel to the fairy tale. They use BlendScape’s saved environments and video positions to replay a history of their collaboration and iterate on their previous scenes to author a new story.
@@
Implementation with BlendScape: We used the inpainting mode to blend the users’ physical environments together. We used image-to-image techniques with prompts of “magic castle, ballroom” and “mushroom forest” to restyle the blended environment to reflect the progression of the story.
Comparison to Prior Work: Existing environment customization tools (e.g., offered by Waazam (Hunter et al., 2014), VideoPlay (Follmer et al., 2010), or Teams TogetherMode) allows users to set the theme by choosing from preset backgrounds or curating and uploading their own images. This results in a manual and time-consuming customization process, but gives end-users fine-grained control over the resulting environment. On the other hand, BlendScape’s image-to-image restyling techniques offer a more rapid and automated method for creating themed environments, which could be more feasible for changing environments to support different collaborative activities during live meetings. BlendScape also enables making granular edits to scenes using GLIGEN inpainting (Li et al., 2023), however, end-users have less control over the output due to the generative model’s stochasticity.
Our review of prior video-conferencing systems (Sec. 2.1) surfaced few examples of using the environment to record a meeting history (e.g., MirrorVerse’s workspace recording tools (Grønbæk et al., 2023a)). BlendScape’s environment history could enable new approaches for facilitating distributed meetings, e.g., capturing visuals that represent discussion topics on the walls of the meeting room and holding later meetings in the same environment to kickstart follow-up conversations.
6. User Study
In the second part of our summative evaluation, we investigated how end-users could envision using BlendScape to personalize video-conferencing environments for various meeting scenarios. We conducted a user study with 15 individuals who regularly use traditional video-conferencing tools (e.g., Microsoft Teams, Zoom), where they used BlendScape to generate video-conferencing environments for various meeting scenarios. Our goals were to (1) Surface end-users’ preferences for customizing meeting environments, including visual or layout characteristics they aimed to achieve; (2) Investigate the benefits and challenges of BlendScape for allowing end-users to express their design intentions.
The study was approved by the Institutional Review Board within our organization. We conducted 1-hour study sessions remotely via Microsoft Teams. Participants were compensated with a $50 gift card.
@@ 6.1. Participants & Recruitment
We recruited individuals from our organization who use video conferencing tools at least three times per week, advertising through internal mailing lists. 15 individuals participated in our study (7 women, 8 men, majority in an age range of 25-34 years) and held a variety of job roles: UX or Product Designer (5), UX Researcher (5), Product Manager (3), Software Engineer (1), and Human Factors Engineer (1). 14 of 15 participants were located in North America and one was located in Asia.
@@ 6.2. Method
The study consisted of three environment composition tasks using BlendScape and a post-task discussion. In the first two tasks, we gradually taught participants to use BlendScape’s composition techniques and asked them to assess the quality of example environments in a semi-structured interview portion. Tasks 1 and 2 served as training for Task 3, where the participants composed a series of video-conferencing environments for a multi-step scenario.
@@
Set-up: We hosted BlendScape on a local machine and gave participants remote control via Microsoft Teams screen-sharing to complete the tasks. For all tasks, we used pre-recorded videos to represent different users (simulated as NDI streams via the OBS Studio NDI Integration Tools888OBS Studio NDI Integration Tools: https://obsproject.com/forum/resources/obs-ndi-newtek-ndi™-integration-into-obs-studio.528/), in order to achieve relatively consistent environment generations across participants. We used a combination of real physical environments and virtual backgrounds to represent a diverse range of environments. For Task 3, we integrated participants’ webcam feeds into BlendScape so they could experience being immersed in different environments.
@@
Task 1: Walkthrough & Comparison of Inpainting Techniques (10 min). First, we explored blended environment designs for a Vacation Planning scenario, where two friends are planning a trip to Paris and generate meeting spaces to reflect landmarks they want to visit. We chose this scenario to teach participants how even basic input (i.e., providing prompts, changing the position and scale of the image priors) can be used to steer the environment generation via BlendScape’s inpainting techniques. After the facilitator demonstrated a few prompts, we asked participants to experiment with different prompts and layouts.
@@
We then showed participants three environments that preserved the users’ webcam backgrounds to varying degrees (Fig. 15). We asked participants to comment on which examples, if any, they could envision using for the Vacation Planning scenario with minimal changes. For the environments they could not envision using, we asked them to explain 1-2 key issues.
@@
Task 2: Walkthrough & Comparison of Image-to-Image Techniques (10 min). Next, we introduced BlendScape’s canvas-based generation techniques using a Game Stream scenario involving a Minecraft player who uses BlendScape to make their viewers feel more engaged and immersed in the gameplay. We demonstrated the image-to-image approach, starting from an image of an arcade and using prompts to restyle the image prior in a Minecraft theme. Then, we demonstrated a more complex approach where we blended the users’ webcam backgrounds into the arcade image. Similarly to the first task, we asked participants to compare four example environments (Appendix A.3), comment on which options they could envision using for the scenario, and explain key issues they observed. We intentionally included examples of roughly blended and cluttered environments to provoke discussions around the limitations of BlendScape’s composition techniques.
@@
Task 3: Designing Environments for a Progressive Meeting Scenario (15 min). To explore how participants would adapt environment designs during a live meeting, we instructed them to compose a series of environments for a scenario involving writing a research paper. First, we introduced a student character (using a pre-recorded video) and brought the participants’ webcam feeds into the prototype, to play the role of another student.
@@
Discussion (15 min). We ended the study with a semi-structured interview around the benefits and limitations that blended representations of environments could provide to remote collaboration scenarios. We asked participants to comment on specific scenarios where they could or would not envision using a system like BlendScape to create personalized and unified environments. Participants also reflected on unexpected or surprising elements of the environments they generated in the previous tasks and how these elements could support or detract from collaborative processes.
@@ 6.3. Data Collection & Analysis
For all study sessions, we collected audio transcripts & recordings, screen recordings, and images of the environments generated with BlendScape for later analysis. To analyze the environment characteristics that participants’ aimed to achieve, we used a thematic analysis approach (Braun and Clarke, 2006). One author reviewed all transcripts to create an initial codebook, specifically paying attention to participants’ rationale for selecting an environment over another in the Task 1 & 2 comparison exercises and revision strategies in Task 3. Three authors then reviewed the codes in the context of specific examples provided by the participants (including both quotes and images of the environments) to extract higher-level themes. We used an affinity diagramming approach (Scupin, 1997) to analyze and aggregate themes from the post-task discussions around the benefits and limitations of BlendScape and future usage scenarios.
@@ 7. Results
In this section, we first present four themes from our analysis of participants’ environment customization preferences (i.e., the visual and layout characteristics participants aimed to achieve in their designs). We then discuss two themes around the benefits of BlendScape’s expressive composition techniques and limitations with the time and effort required to achieve optimal designs. Figure 12 shows a sample of the environments generated by the 15 participants in our study.
@@ 7.1. Environment Customization Preferences
We discuss four themes around participants’ preferences for (1) authentic over artificial spaces; (2) strongly vs. subtly themed environments; (3) structuring collaboration through spatial layouts of users; (4) balancing the spatial-richness of environments.
@@
A majority of participants expressed that blending users’ physical surroundings into a unified meeting space can promote co-presence while maintaining familiar aspects of their individual environments (P2, P5-6, P9, P11-14). P13 commented that “taking something personal to [them] and tweaking it” would help remote collaborators to “trick [their] minds to believe that [they’re] more in the same place.”
P6 cited another benefit of authentic backgrounds in creating a higher degree of realism, as users’ videos appeared to be naturally framed with appropriate furniture and lighting. In environments constructed from virtual backgrounds, users sometimes appeared to be “floating in space in front of an image… like a bad Photoshop job” (P6). However, we observed a drawback to this increased sense of realism: participants more easily identified and were more critical of flaws in how their physical surroundings were blended (P1, P3-5, P7-8, P10), e.g., warped areas or inconsistent room geometry (Fig. 13). With fully artificial environments, some participants argued that “there’s no pressure to do it well” (P8); “even if it was not as polished” they would have a “higher tolerance” for mistakes (P3).
@@
A few participants appreciated having stylistic elements that closely reflect the Meeting Activity and Theme prompts they provided, e.g., a chemistry lab for the Research Paper scenario in Task 3. Despite some unrealistic visuals, P15 appreciated the departure from traditional meeting spaces that are “not really imaginative,” “fixed and static”. P7 commented: “it’s nice that we don’t need the space to look like traditional meeting spaces.” However, most participants preferred simple and subtle theming to avoid distractions (P1, P7, P9, P13-15); P1 expressed environments should “add texture to the call without pulling away from it.”
Environments that establish spatial layouts of users can help structure collaboration, but require manual staging.
A popular design strategy was placing users around spatial landmarks in the scene to reflect the collaboration context (P1, P3-7, P9, P13-15). For example, in the Research Paper scenario, P14 positioned the professor in the lower left corner “to help structure the collaboration,” similar to a picture-in-picture style for online lectures. P9 placed users in chairs that were close together to “give [them] the most privacy” while having a one-on-one discussion.
However, some participants observed that BlendScape did not always consider the existing layout of users when generating meeting spaces. P1 and P6 had to manually reposition users to seat them in chairs, but they expected BlendScape to automatically generate chairs and tables to frame the users. This points to a limitation of how BlendScape blends environments after segmenting users from their video backgrounds, which we further discuss in Sec. 8.2.
Balancing the spatial richness of environments with 2D representations of users.
Spatial properties of BlendScape’s environments (e.g., furniture layouts, lighting, and shadows aligned with the geometry of the space) increased realism and lent themselves to structuring user layouts for collaboration (as discussed in the previous theme). However, many participants observed a juxtaposition between these spatially-rich environments and the strictly 2-dimensional representation of users (P5-6, P8-10, P11, P13-14). A few participants tried to tilt the users’ videos to “echo the [3D] perspective in the room” (P14) and make it look like users were facing each other (P9, P13). To make our hidden surface removal technique (i.e., hiding floating heads behind foreground objects) more convincing, P14 suggested to “accept the flatness” and use environments where all furniture and users’ videos are facing forwards, in the style of Teams TogetherMode.

### v2

@@ 5. Demonstration of Additional Scenarios
In this section, we illustrate the expressiveness of BlendScape’s generative AI-enabled customization tools through prototyping meeting environments for three distributed collaboration scenarios: Design Brainstorming, Remote Education, and Storytime with Family.
Earlier, we reviewed the range of video-conferencing environments implemented by prior systems to enhance distributed collaborators’ communication and sense of co-presence (Sec. 2.1). This review surfaced ten environment design strategies, such as incorporating representations of task spaces, enabling proxemic metaphors, or recording a history of collaborative activities. We demonstrate how BlendScape’s composition techniques can be used to implement eight out of the ten design strategies (shown in bold in Fig. 2).
Our focus for BlendScape was enabling end-users to compose custom environment visuals; as such, rendering spatial sound was out of scope (Fig. 2B.iii). Currently, BlendScape’s 2D image generation techniques do not support rendering egocentric viewing perspectives to simulate face-to-face communication cues (Fig. 2B.v), e.g., turning to face other users seated around a table (Tang et al., 2023). This could be achieved via 360∘ panoramic or 3D environment generation approaches (Numan et al., 2024; Li and Bansal, 2023).
@@ 5.1. Design Brainstorming
Recall our initial scenario from Sec. 3, where two interaction designers are remotely collaborating to prototype mixed reality interfaces (Fig. 3). They use BlendScape to render their videos in a unified workspace that blends their physical and digital task spaces into a single design studio. This allows them preview their mixed reality designs and simulate face-to-face design critiques, by verbally referring to and gesturing around hand-drawn sketches.
Implementation with BlendScape: We prototyped this scenario using four live video feeds: two webcam feeds of the designers, an external camera feed capturing a physical desk with notebooks, and a screen-capture of a tablet-based digital sketching application (Fig. 3a). We used BlendScape’s inpainting mode to blend the users’ videos and the task space feeds into seamless meeting environments (Fig. 3b, c). Finding the ideal placement of the task space feeds required some trial-and-error; for example, positioning the physical sketches at the bottom of the screen occasionally rendered them on the floor of the environment rather than on a desk.
@@ 5.2. Remote Education
Next, to demonstrate how BlendScape environments could support spatial metaphors to facilitate conversations (Hu et al., 2023a), we prototyped a scenario involving a remote university course.
Scenario Description: A professor uses BlendScape to facilitate activities in their literature seminar. Before the seminar starts, they establish the room layout for small group discussions: they upload an image of a room with several tables and restyle it as a “library” (Fig. 10a, c). Students assign themselves to groups by organizing around the tables, leveraging proxemic metaphors (Fig. 10c). To structure a large group discussion later in the seminar, the professor restyles an image of a conference room (Fig. 10b, d). To spotlight speakers, they use scaling metaphors by rendering the students’ videos larger and at the head of the table.
Implementation with BlendScape: We selected two images priors of event spaces that contained enough seating for 5 users and were captured from similar forward-facing perspectives as the users’ webcam videos, allowing us to realistically place users behind furniture. We then used BlendScape’s image-to-image mode with a Meeting Theme prompt of “library” to restyle the image priors.
@@ 5.3. Storytime with Family
Scenario Description: A grandmother uses BlendScape to craft a memorable storytelling experience with her granddaughter. At the start of the call, a castle appears in the space between their video backgrounds with a prompt of “storytelling” (Fig. 11b). As the grandmother reads a fairy tale, she restyles the scenes to set the theme of the meeting; the environment transforms into magic ballroom and then a mushroom forest (Fig. 11c, e). These environments illustrating the story’s progression serve as an artifact produced through their collaboration. A few weeks later, the granddaughter wants to write a sequel to the fairy tale. They use BlendScape’s saved environments and video positions to replay a history of their collaboration and extend their previous scenes.
Implementation with BlendScape: We used the inpainting mode to blend the users’ physical environments together. We used image-to-image techniques with prompts of “magic castle, ballroom” and “mushroom forest” to restyle the blended environments to reflect the events of the story.
6. Exploratory Study with End-Users
To investigate whether and how end-users would find value in using generative AI to personalize video-conferencing environments, we conducted an exploratory study with 15 frequent users of traditional video-conferencing tools (e.g., Microsoft Teams, Zoom). Our goals were to (1) surface participants’ preferences for customizing meeting environments, including visual or layout characteristics they aimed to achieve; (2) investigate to what extent BlendScape allowed participants to express their design intentions.
The study was approved under the Institutional Review Board (Ethics / IRB ID: #10764; Release and Compliance ID: #6755). We conducted 1-hour study sessions remotely via Microsoft Teams. Participants were compensated with $50 gift cards.
@@ 6.1. Participants & Recruitment
Advertising through internal mailing lists, we recruited individuals from our organization who use video conferencing tools at least three times per week. 15 individuals participated in our study (7 women, 8 men, majority in an age range of 25-34 years) and held a variety of job roles: UX or Product Designer (5), UX Researcher (5), Product Manager (3), Software Engineer (1), and Human Factors Engineer (1). 7 out of 15 participants used image generation models once a week or more frequently; the remaining 8 used them once a month or less frequently. 14 of 15 participants were located in North America and one was located in Asia.
@@ 6.2. Method
The study consisted of three environment composition tasks using BlendScape and a post-task discussion. In the first two tasks, we introduced participants to BlendScape’s inpainting and image-to-image techniques, teaching them to generate environments based on pre-defined prompts and assess the quality in a semi-structured interview portion. This served as training for Task 3, where participants combined BlendScape’s composition tools to create a series of environments for a research meeting scenario.
@@
Set-up: We hosted BlendScape on a local machine and gave participants remote control via Microsoft Teams screen-sharing. For all tasks, we used pre-recorded videos to represent different users (simulated as NDI streams via the OBS Studio NDI Integration Tools666OBS Studio NDI Integration Tools: https://obsproject.com/forum/resources/obs-ndi-newtek-ndi™-integration-into-obs-studio.528/), in order to achieve relatively consistent environment generations across participants. We used a combination of real physical locations and virtual backgrounds to represent a diverse range of image priors. For Task 3, we integrated participants’ webcam feeds into BlendScape so they could experience being immersed in different environments.
@@
Task 1: Walkthrough & Comparison of Inpainting Techniques (10 min). First, we explored blended environment designs for a Vacation Planning scenario, where two friends are planning a trip to Paris and generate meeting spaces to reflect landmarks they want to visit. The facilitator illustrated how even basic input (i.e., providing prompts, changing the position and scale of the image priors) can be used to steer the environment generation via BlendScape’s inpainting techniques. Afterwards, we asked participants to experiment with different prompts and layouts.
@@
We then showed participants three environments that preserved varying degrees of users’ video backgrounds (Fig. 15). We asked participants to comment on which examples, if any, they could envision using for the Vacation Planning scenario with minimal changes. For the environments they could not envision using, we asked them to explain 1-2 key issues.
@@
Task 2: Walkthrough & Comparison of Image-to-Image Techniques (10 min). Next, we introduced BlendScape’s image-to-image generation techniques using a Game Stream scenario, where a Minecraft player uses BlendScape to engage viewers their viewers. We started from an image of an arcade and used prompts to restyle the image prior in a Minecraft theme. Then, we combined inpainting and image-to-image techniques to blend the users’ webcam backgrounds into the arcade image. Similarly to the first task, we asked participants to compare four example environments (Fig. 16), comment on which options they could envision using for the scenario, and explain key issues they observed. We intentionally included examples of roughly blended and cluttered environments to provoke discussions around the limitations of BlendScape’s composition techniques.
@@
Task 3: Designing Environments for a Progressive Meeting Scenario (15 min). To explore how participants could envision adapting environment designs during a live meeting, we instructed them to compose a series of scenes for a scenario involving writing a research paper. First, we introduced a student character (using a pre-recorded video) and brought the participants’ webcam feeds into BlendScape, to play the role of another student.
@@
Discussion (15 min). We ended with a semi-structured interview around the potential value that generative AI-enabled custom environments could provide to distributed collaboration scenarios. We asked participants to comment on specific scenarios where they could or could not envision using a system like BlendScape to personalize meeting environments. Participants also reflected on unexpected or surprising elements of the environments they generated in the previous tasks and how these aspects might support or detract from collaborative processes.
@@ 6.3. Data Collection & Analysis
For all study sessions, we collected audio transcripts & recordings, screen recordings, and images of the environments generated with BlendScape for later analysis. To analyze the environment characteristics that participants aimed to achieve, we used a thematic analysis approach (Braun and Clarke, 2006). One author reviewed all transcripts to create an initial codebook, specifically paying attention to participants’ rationale for selecting an environment over another in the Task 1 & 2 comparison exercises and their revision strategies in Task 3. Two authors then reviewed the codes in the context of specific examples provided by the participants (including both quotes and images of the environments) to extract higher-level themes. We used an affinity diagramming approach (Scupin, 1997) to analyze and aggregate themes from the post-task discussions around the benefits and limitations of BlendScape and future usage scenarios.
@@ 7. Results
Figure 12 shows a sample of the environments generated by the 15 participants in our study. In this section, we first present four themes around participants’ environment customization preferences (i.e., the visual and layout characteristics they aimed to achieve). We then discuss the benefits of BlendScape’s composition techniques for expressive leverage and limitations with the time and effort required to achieve optimal designs.
@@ 7.1. Environment Customization Preferences
When assessing the quality of BlendScape’s environments, participants expressed preferences for (1) authentic over artificial spaces; (2) both strong and subtle thematic elements, depending on the meeting context; (3) structuring collaboration through spatial layouts of users; (4) balancing the spatial-richness of environments.
@@
A majority of participants expressed that blending users’ physical surroundings into a unified meeting space could promote co-presence while maintaining familiar aspects of their individual environments (P2, P5-6, P9, P11-14). P13 commented that “taking something personal to [them] and tweaking it” would help remote collaborators to “trick [their] minds to believe that [they’re] more in the same place.”
P6 cited another benefit of authentic backgrounds in creating a higher degree of realism, as users’ videos appeared to be naturally framed with appropriate furniture and lighting. In environments constructed from virtual backgrounds, users sometimes appeared to be “floating in space in front of an image… like a bad Photoshop job” (P6). However, we observed a drawback to this increased realism: participants more easily identified and were more critical of flaws in how their own physical surroundings were blended (P1, P3-5, P7-8, P10), e.g., warped areas or inconsistent room geometry (Fig. 13). With fully artificial environments, some participants argued that “there’s no pressure to do it well” (P8); “even if it was not as polished” they would have a “higher tolerance” for mistakes (P3).
@@
A few participants appreciated having stylistic elements that closely reflect the Meeting Activity and Theme prompts they provided, e.g., a chemistry lab for the Research Paper scenario in Task 3. These participants embraced the at-times unrealistic visuals, arguing “that we don’t need the space to look like traditional meeting spaces” (P7) which are “fixed and static… not really imaginative” (P15). However, most participants preferred simple and subtle theming to avoid distractions (P1, P7, P9, P13-15). As P1 stated, environments should “add texture to the call without pulling away from it.”
Establishing user layouts through the environment was a popular design strategy, but required manual staging.
Many participants placed users around spatial landmarks in the scene to reflect specific collaborative activities (P1, P3-7, P9, P13-15). For example, in the Research Paper scenario, P14 positioned the professor in the lower left corner “to help structure the collaboration,” similar to a picture-in-picture style for online lectures. P9 placed users in chairs that were close together to “give [them] the most privacy” while having a one-on-one discussion.
However, some participants observed that BlendScape did not generate spaces to accommodate the existing layout of users in the scene. P1 and P6 had to manually reposition users to seat them in chairs, but they expected BlendScape to automatically generate chairs and tables to frame the users. This points to a limitation of how BlendScape blends environments after segmenting users from their video backgrounds, which we further discuss in Sec. 8.2.
Need to balance the spatial richness of environments with 2D representations of users.
Participants found that spatial properties of BlendScape’s environments (e.g., furniture layouts and lighting aligned with the geometry of the space) increased realism and lent themselves to structuring user layouts for collaboration (as discussed in the previous theme). However, many participants observed a juxtaposition between these spatially-rich environments and the strictly 2D representation of users (P5-6, P8-10, P11, P13-14). A few participants tried to tilt the users’ videos to “echo the [3D] perspective in the room” (P14) and make it look like users were facing each other (P9, P13). To make our hidden surface removal technique (i.e., hiding floating heads behind foreground objects) more convincing, P14 suggested to “accept the flatness” and use environments where all furniture and users’ videos are facing forwards, in the style of Teams TogetherMode11footnotemark: 1.